OpenStudy (anonymous):

Also number 19 please! 18 is 7.143

OpenStudy (anonymous):

i am extremely stuck on this one

OpenStudy (anonymous):

maximum height = 90 and maximum range = double of half range = 12 90 =u^2sinthita^2/2g 12 =u^2sin2thita/g divide the 2 to get 90/12 = tanthita/4 tan thita = 30 since tan thita = 30 sin thita = 30/rooot(901) put this in 12 =u^2sin2thita/g and get the value of g

OpenStudy (anonymous):

wait what is u^2??